OpenStudy (anonymous):

Solve the equation: 4x2 = 9 A. {-2, -3} B. {2/3, -2/3} C. {3/2, -3/2} D. no solution

OpenStudy (anonymous):

x^2 = 9/4 = (3/2)2 x = 3/2, -3/2 D.